# Kuznecov formulae for fractal measures

### v1

@@ 5.2. Necessity of the averaged $s$-density condition
The averaged $s$-density hypothesis in Theorem 1.3 is not merely a technical convenience: without a stabilization of the small-scale behavior of $\mu$, one cannot in general expect the normalized quantities $N_{\mu}(\lambda)/\lambda^{n-s}$ to converge.
@@
If $\mu$ is $s$-Ahlfors regular then $a(r)$ is bounded above and below by positive constants for $0<r\leq r_{0}$.
Since $F(0)=0$, integration by parts gives, modulo a boundary term that is exponentially decaying in $1/t$, we have
@@
Lemma 5.1 (Realizing $\liminf/\limsup$ from symmetric scale blocks).
@@ 5.3. Optimality of the remainder
In the smooth submanifold case, Zelditch’s Kuznecov formulae (1.2) come with an $O(\lambda^{n-k-1})$ remainder term. Moreover, for generic metrics this remainder can be improved to $o(\lambda^{n-k-1})$ by a recent result of Kaloshin, Wyman and the author [KWX25]. In contrast, the asymptotic formula in Theorem 1.3 is stated with an $o(\lambda^{n-s})$ remainder. It is natural to ask whether, under the assumptions of $s$-Ahlfors regularity and the existence of an averaged $s$-density constant, one can improve this to a uniform power-saving error term of the form
@@
for some $\delta>0$ depending only on $(M,g,n,s)$ and valid for all measures $\mu$ in our class. We now explain why such a uniform improvement is not available in general.
@@
Then $\tfrac{3}{4}\leq w\leq\tfrac{5}{4}$, so $\mu$ is $s$-Ahlfors regular. Moreover, since $\mu$ is absolutely continuous with respect to $dV_{H}$ with bounded density $w$, $\mu$ admits a pointwise $s$-density $\theta=w$. Hence, by Remark 1.4,

### v2

@@ 5.1. Necessity of the averaged $s$-density condition
In this subsection we prove Proposition 1.7, which shows that the averaged $s$-density hypothesis in Theorem 1.3 is not merely a technical convenience. Without a stabilization of the small-scale behavior of $\mu$, one cannot in general expect the normalized quantities $N_{\mu}(\lambda)/\lambda^{n-s}$ to converge.
@@
If $\mu$ is $s$-Ahlfors–David regular then $a(r)$ is bounded above and below by positive constants for $0<r\leq r_{0}$.
Since $F(0)=0$, an integration by parts yields, up to a boundary term that decays exponentially in $1/t$,
@@
Lemma 5.1 (Realizing $\liminf/\limsup$).
@@ 5.2. Optimality of the remainder
In the smooth submanifold case, Zelditch’s Kuznecov formula (1.2) comes with an $O(\lambda^{n-k-1})$ remainder term. Moreover, for generic metrics this remainder can be improved to $o(\lambda^{n-k-1})$ by a recent result of Kaloshin, Wyman and the author [KWX25]. In contrast, the asymptotic formula in Theorem 1.3 is stated with an $o(\lambda^{n-s})$ remainder.
It is natural to ask whether one can strengthen (1.7) to a uniform power-saving error term of the form
@@
for some $\delta>0$ depending only on $(M,g,n,s)$ and valid for all measures $\mu$ satisfying the hypotheses of Theorem 1.3. Proposition 1.8 asserts that such a uniform improvement is not available in general, even under additional regularity assumptions, and we prove it in this subsection.
@@
Then $\tfrac{3}{4}\leq w\leq\tfrac{5}{4}$, so $\mu$ is $s$-Ahlfors–David regular. Moreover, since $\mu$ is absolutely continuous with respect to $dV_{H}$ with bounded density $w$, $\mu$ admits a pointwise $s$-density $\theta=w$. Hence, by Remark 1.5,
@@
5.3. A self-similar example on the unit interval
We record an explicit self-similar fractal measure on $[0,1]$ of non-integer dimension satisfying the averaged $s$-density condition (1.5).
Proposition 5.3.
Let $\phi_{0},\phi_{1}:[0,1]\to[0,1]$ be
$\phi_{0}(x)=\tfrac{1}{2}x,\qquad\phi_{1}(x)=\tfrac{2}{3}+\tfrac{1}{3}x,$
and let $K\subset[0,1]$ be the attractor $K=\phi_{0}(K)\cup\phi_{1}(K)$. Let $s\in(0,1)$ be the unique solution of $2^{-s}+3^{-s}=1$, and set $p_{0}:=2^{-s}$, $p_{1}:=3^{-s}$. Let $\mu$ be the unique Borel probability measure supported on $K$ satisfying
(5.16) $\mu(A)=p_{0}\,\mu(\phi_{0}^{-1}(A))+p_{1}\,\mu(\phi_{1}^{-1}(A))\qquad\text{for all Borel }A\subset[0,1].$
Then $s\notin\mathbb{N}$, and $\mu$ admits an averaged $s$-density constant $A_{\mu}\in(0,\infty)$ in the sense of (1.5).
Proof.
It is clear that $s$ is not an integer. For $0<r<1/6$, pairs from different first-level pieces cannot satisfy $|x-y|\leq r$, hence expanding $\mu\times\mu$ using (5.16) gives the exact recursion
$F(r)=p_{0}^{2}\,F(2r)+p_{1}^{2}\,F(3r)\qquad(0<r<1/6).$
Setting $f(r):=F(r)/(\operatorname{vol}(B^{s})\,r^{s})$ and using $p_{0}=2^{-s}$, $p_{1}=3^{-s}$ yields
$f(r)=p_{0}\,f(2r)+p_{1}\,f(3r)\qquad(0<r<1/6).$
With $t=-\log r$ and $g(t):=f(e^{-t})$, we obtain
$g(t)=p_{0}\,g(t-\log 2)+p_{1}\,g(t-\log 3)\qquad(t>\log 6).$
Since $\log 2/\log 3\notin\mathbb{Q}$, the key renewal theorem (see, e.g. [Fel71]) implies that $g(t)\to A_{\mu}$ as $t\to\infty$ for some $A_{\mu}\in(0,\infty)$, i.e. $f(r)\to A_{\mu}$ as $r\to 0^{+}$, which is exactly (1.5) with exponent $s$. Finally, we remark that this measure $\mu$ in our example satisfies Ahlfors–David regularity. ∎